OpenStudy (anonymous):

please help! a sail is in the form of a right triangle that is 3 times as high as it is wide. the sail is made from 6 meters of material. what is the height? remember a=1/2 bh

OpenStudy (anonymous):

6m = h + w 6 m = 3w + w 4w= 6 w= 1.5 so, height = h= 3w = 3*1.5 = 4.5 m

OpenStudy (anonymous):

h = 3w A = 1/2 3 w w = 3/2 w^2 Is it 6 m^2 of material? If so we have 2/3 times 6 = 4 = w^2 w = 2 h = 6